# Gravity versus astrophysics in black hole images and photon rings: equatorial emissions and spherically symmetric spacetimes

### v1

@@ 2.3 Apparent shape of circular rings
Figure 2: Direct (solid lines in the upper panels) and first-lensed (dashed lines in the lower panels) apparent positions of rays emitted from isoradial distances $r$ to the black hole at different inclinations: $r=2M$ (grey), $2.5M$ (magenta), $3M$ (green) or $5.5M$ (yellow). The inner shadow (brown) and the first lensing band (light blue) are represented as well.
@@
We analyse projections of equatorial rings of constant radii $r$. At this point we are only interested in mapping between a given equatorial plane and the observer’s screen, connected with null geodesics. Hence, we do not need to specify neither emission and opacity properties nor redshift effects. Furthermore, we restrain our attention to light rays of orders $n=0,1$, but the following enquiries can be generalised to higher order. Discussions presented herein are illustrated in Figure 2.
@@
If the equatorial plane is observed face-on, the projections of a ring of radius $r$ within the disc are circles on the plane of sky. The arrival impact parameter $b$ of a photon emitted from a radius $r>r_{\mathrm{H}}$ only depends on $r$ and can be approximated by ‘just adding one’ (Gates et al. 2020). This relation, valid for a Kerr black hole, can be rewritten for a Schwarzschild black hole as
[…]
@@
Hence, the event horizon, situated at the radius $r=2M$, will appear at an apparent approximated radius $b\approx 2.82M$, in agreement with the numerical outputs of Figure 2, while outer radii develop towards higher values of $b$. As for geodesics of order $n=1$, if one stops the backward ray-tracing integration, properly introduced in Section 5, before getting the third crossing of the plane, the numerical inner and outer edges of the $n=1$ lensing band will appear at $b=5.02M$ and $b=6.17M$ respectively (Gralla et al. 2019); these values agree with Figure 2.
@@
If the screen and the equatorial plane are not parallel, the images of equatorial rings are deformed by projection effects [^8] In a Newtonian context, they would give rise to ellipses.. On the contrary, the critical curve does not vary under change of the plane’s inclination $i$ because of the spherical symmetry of the photon sphere and it always takes the shape of a circle of radius $b=3\sqrt{3}M$ on the plane of sky [^9] In Kerr, the photon shell has an equatorial reflection symmetry and the critical curve is always symmetric with respect to the axis perpendicular to the projected direction of the spin (Gralla & Lupsasca 2020a)..
@@
Let us focus now on the general properties of projections on the observer’s screen of equatorial circles as seen by an inclined observer. The discussions of the following paragraphs are complemented by additional details given in Appendix A. First of all, both the zeroth-order and the first-lensed projections of planar circular rings are equal to their non-inclined counterparts along the line on the screen parallel to the inclination’s axis of rotation of the plane (horizontal axis in the upper panels of Figure 1). Moreover, projected rings always possess a mirror symmetry on the observer’s screen with respect to the axis that is normal to the inclination’s one [^10] This property, reliant on the planar motion typical of all static spherical symmetric spacetimes, is broken in Kerr (Gralla & Lupsasca 2020a). (vertical axis in Figure 1).
@@
Let $\psi$ be the polar angle on the observer’s screen as shown in Figure 1. If the radius of the inclined ring satisfies $r_{\mathrm{H}}<r\leq\tilde{r}$, the zeroth-order projection of the side of the ring tilted backward ($0^{\circ}<\psi<180^{\circ}$, ) or forward ($180^{\circ}<\psi<360^{\circ}$) appears, respectively, at a higher or lower radial distances $b$ on the screen with respect to the non-inclined apparent circle. As a result, since in our simulations the forward side of the plane points towards $\psi=270^{\circ}$, the impact parameter $b$ of the inclined projected equatorial horizon, as well as that of rings with $r\leq 3M$, decreases from $\psi=90^{\circ}$ to $\psi=270^{\circ}$, see grey and magenta lines in the upper panels of Figure 2. The same behaviour of $b$ is followed for the forward part of rings with $r>3M$: it then decreases from $\psi=180^{\circ}$ to $\psi=270^{\circ}$ in our set-up, see green and yellow lines in the upper panels of Figure 2. Instead, for the backward points of rings with $r>3M$, for a given pair of plane’s inclination and direction on the observer’s screen, there exist a radius for which the behaviour of $b$ is reversed, that is to say that the projection of a given point has a smaller impact parameter than its corresponding non-inclined projection. The existence of this radius, which is smaller for higher inclinations and for directions on the screen closer to the axis of reflection, is a necessary consequence of the flat spacetime limit: in a Minkowski spacetime, the projected image of a circle is an ellipse whose semi-major axis is along the rotation axis (the horizontal direction here) and equal to the non-inclined projection. Since the semi-minor axis is smaller than the semi-major axis by definition, a reversal must appear at large distance from the black hole to reproduce the Minkowski behaviour. For the small inclination considered in the set-up of Figure 2, this effect appears beyond the considered field of view.
@@
As for the first-lensed projection of equatorial rings, if the radius of the ring satisfies $r_{\mathrm{H}}<r\leq\tilde{r}$, the projection of the forward or backward side of the ring appears, respectively, at a higher or lower radial distances $b$ on the screen with respect to the non-inclined apparent circle. Hence, as the geodesic executes a half turn around the black hole before touching the screen and since in our simulations the forward side of the plane points towards $\psi=270^{\circ}$, the impact parameter $b$ of such rings decreases from $\psi=90^{\circ}$ to $\psi=270^{\circ}$, see grey and magenta lines in the lower panels of Figure 2. If $r=3M$, there is no variation with respect to the non-inclined case, see green line in the lower panels of Figure 2. Finally, if $r>3M$, the modulations of $b$ are reversed. Thus, in our simulations, $b$ increases from $\psi=90^{\circ}$ to $\psi=270^{\circ}$, see yellow line in the lower panels of Figure 2. Put differently, the ring projections flee the critical curve in the lower part of the screen and are attracted to it in the upper part.
2.4 Inner shadow, primary image and photon rings
@@
The relevant observables appear on the far-away observer’s screen if and only if emitting material is present around the black hole. In particular, a single point source gives rise to an infinite sequence of images. These additional lensed observables were originally called ‘ghost images’ (Darwin 1959) and were present in the first simulated ‘photograph’ of an accretion disc around a Schwarzschild [^11] The first image of an accretion disc around a Kerr black hole with $a\neq 0$ has been computed by Viergutz (Viergutz 1993). black hole (Luminet 1979). SMBHs like Sgr A* and M87* are surrounded by emitting matter forming a hot, geometrically thick and optically thin radiatively inefficient accretion flow (RIAF, Yuan & Narayan 2014). Moreover, for M87*, EHT data favours the scenario of a magnetically arrested disc (MAD) whose millimetre wavelength emission arises near the event horizon, predominantly in the vicinity of the equatorial plane (Narayan et al. 2003; The Event Horizon Telescope Collaboration 2019c; The Event Horizon Telescope Collaboration 2021).
@@
Images of black holes surrounded by optically thin emission display a brightness depression in their inner region, known as observable shadow, that does not always track the critical curve, but it is instead highly astrophysical dependent (Chael et al. 2021). For instance, for an idealistically thin equatorial disc, this ‘inner shadow’ can be well inside the critical curve: when the disc extends below the photon sphere, the edge of this darkest region corresponds to the projection of the inner radius of the equatorial disc modulated by redshift effects (Chael et al. 2021). For a disc seen face-on and extending to the event horizon of a Schwarzschild black hole, the inner shadow has a radius of about $2.82M$ on the observer’s screen, as discussed in Paragraph 2.3.
@@
The primary image is the one reaching the observer on geodesics touching the emitting equatorial source just once. It is also called $n=0$ or direct image. Photon rings are higher order images of the same emitting material, labelled by the number $n\geq 1$ of the equatorial-plane crossings along their nearly bound orbit. When the source is intersected during the crossing, a flux increment is added. Also, the $n$-th photon ring of the equatorial source lies inside the $n$-th lensing band (Gralla & Lupsasca 2020a).
@@
When the emitting source is not isotropically distributed, as in the case of equatorial discs, photon rings exponentially tend to the critical curve, have exponentially decreasing angular width and thus are exponentially demagnified and exponentially loose flux density [^12] The first M87* photon ring provides $\sim 10\%$ of the total luminosity in numerical GRMHD simulations (Johnson et al. 2020). for increasing $n$ (Johnson et al. 2020; Vincent et al. 2022). Also, in intensity cross sections of the images, they do not blend with the direct image, unlike an infalling sphere, but they give birth to a stack of discrete higher order images superimposed on top of the primary one (Gralla et al. 2019; Vincent et al. 2022), provided that absorption effects are weak at a given observing frequency (Beckwith & Done 2005). These properties are presented more meticulously in the introduction of Paugnat et al. (2022).
@@
To summarise, photon rings are signatures related to the strong-field lensing effects of the photon sphere and, unlike the critical curve or the lensing bands, they can be observed and possess only limited astrophysical dependence (Vincent et al. 2022). The present EHT does not resolve the photon rings (Gralla 2021), but their interferometric detection may soon become possible either by observing at higher frequencies (ngEHT, Johnson et al. 2023) or on longer Earth-space baselines (BHEX, Johnson et al. 2024).
@@ 3.1 M87* distance and orientation
Our compact object model mimics the observationally assessed properties of M87*, namely its mass $M$, its distance $D$ and its inclination $i$. The values of $M$ and $D$ are those inferred by the EHT collaboration (The Event Horizon Telescope Collaboration 2019d), the mass being consistent with that obtained via a stellar-dynamics study (Gebhardt et al. 2011). For its inclination, defined as the angle between the line of sight and the angular momentum of the accretion flow, we use the value of the viewing angle of the jet in M87 (Mertens et al. 2016), assuming that its axis is aligned with the angular momentum vector of the disc and that the latter is pointing away from the observer.
@@ 3.2 Rezzolla-Zhidenko metric
In this paper, we consider the Rezzolla-Zhidenko metric (RZ metric, Rezzolla & Zhidenko 2014) which describes, in a parametric framework, the spacetime outside a general spherically symmetric [^13] The Konoplya-Rezzolla-Zhidenko metric (KRZ metric, Konoplya et al. 2016) generalises the RZ parametrisation to circular spacetimes, then to rotating black holes, but not all stationary axisymmetric black holes are included in this class. stationary, and therefore static, black hole [^14] For an extension of the RZ scheme to arbitrary asymptotically flat, spherical symmetric and static spacetimes, including non-BH ones like boson stars or naked singularities, see Kocherlakota & Rezzolla (2020).. In the strong-field regime, contrary to previous parametrisations in which an infinite number of equally important parameters is needed (Johannsen & Psaltis 2011), the RZ metric is described by a finite number of hierarchical factors. The line element of this metric in a spherical polar coordinate system $(t,r,\theta,\varphi)$ is:
@@
where the functions $N$ and $B$ only depend on the radial coordinate $r$. Conveniently for a black hole, $r$ can be compactified in the dimensionless variable $x\coloneqq 1-r_{\mathrm{H}}/r$, where $r_{\mathrm{H}}$ is the radial position of the event horizon, so that $x=0$ corresponds to the location of the horizon and $x=1$ to spatial infinity. In terms of $x$, using $G=c=1$ natural units and introducing $\epsilon\coloneqq 2M/r_{\mathrm{H}}-1$, the functions $N$ and $B$ can be written as follows:
[…]
@@
$\tilde{A}(x)=\frac{a_{1}}{1+\dfrac{a_{2}\,x}{1+\dfrac{a_{3}\,x}{1+\ldots}}}\quad\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},\quad\tilde{B}(x)=\frac{b_{1}}{1+\dfrac{b_{2}\,x}{1+\dfrac{b_{3}\,x}{1+\ldots}}}\>\>\,.$ (6)
The dimensionless constants $a_{0},a_{1},a_{2},a_{3}\ldots$ and $b_{0},b_{1},b_{2},b_{3}\ldots$ have to be constrained by observations. If all $\epsilon$, $a_{i}$ and $b_{i}$ are zero, with $i\in\mathbb{N}$, the Schwarzschild metric is recovered in Schwarzschild-Droste coordinates.
It is useful to notice that at the horizon the functions $\tilde{A}$, $\tilde{B}$ of equations (6), reduce to $\tilde{A}(0)=a_{1}$, $\tilde{B}(0)=b_{1}$. This means that near the horizon, only the lower-order terms in the expansion are relevant. Higher-order terms increase the accuracy of the approximation and, for numerous known metrics deviating slightly from the Schwarzschild one, the first order is sufficient to accurately investigate astrophysical observable deviations (Konoplya & Zhidenko 2020). Besides, the coefficients $a_{0}$ and $b_{0}$, that have been introduced in the equations (5), can be expressed as a combination of the parametrised post-Newtonian (PPN) parameters and, assuming black hole uniqueness, they are constrained by observations to be at most of the order of $10^{-4}$ (Will 2014).
For the sake of simplicity, as we are interested in the near-horizon emission around a black hole and following the arguments stated before, in the simulations all the parameters $a_{i}$, $b_{i}$ with $i\geq 2$ are put to zero. The parameters $a_{0}$ and $b_{0}$ are also fixed to be zero, their impact on the image formation being negligible with respect to that of the astrophysical set-up when PPN constraints are met, as it is illustrated in Appendix C; in the same Appendix we also present higher deviations in $a_{0}$ and $b_{0}$, allowed in theories of gravity without a Birkhoff-like uniqueness theorem. Finally, we choose $b_{1}=0$, again for simplicity purposes and because its variation has a very limited effect on the image, as shown in Section 6 and Appendix C. Then, we can study the effect on the images of the only unconstrained parameters $a_{1}$ and $\epsilon$ that we will make vary independently, specifically fixing the other one to zero, in their theoretically-allowed [^15] The outermost horizon is located at $r=r_{\mathrm{H}}$ and no larger roots of $g_{rr}^{-1}=N^{2}(r)$ should exist (Kocherlakota & Rezzolla 2022). ranges and taking ‘small’, in the sense of the following right inequality [^16] This does not imply that the deviations are physically small., parametric deviations (Kocherlakota & Rezzolla 2020):
$-1<\epsilon\leq 0.5,\quad a_{1}\geq-1\quad\text{and}\quad\max{(\,|\,a_{1}\,|,|\,\epsilon\,|\,)}\leq 1\>.$
Table 1: Parameters of the black hole

### v2

@@ 2.3 Apparent shape of circular rings
Figure 2: Direct (solid lines in the upper panels) and first-lensed (dashed lines in the lower panels) apparent positions of rays emitted from isoradial distances, $r$, to the black hole at different inclinations: $r=2M$ (grey), $2.5M$ (magenta), $3M$ (green), and $5.5M$ (yellow). The inner shadow (brown) and the first lensing band (light blue) are represented as well.
@@
We analysed the projections of equatorial rings of constant radii, $r$. At this point we were only interested in mapping between a given equatorial plane and the observer’s screen, connected with null geodesics. Hence, we did not need to specify the emission and opacity properties nor the redshift effects. Furthermore, we restrained our attention to light rays of orders $n=0,1$, but the following enquiries can be generalised to a higher order. Discussions presented herein are illustrated in Figure 2.
@@
If the equatorial plane is observed face-on, the projections of a ring of radius $r$ within the disc are circles on the plane of sky. The arrival impact parameter, $b$, of a photon emitted from a radius $r>r_{\mathrm{H}}$ only depends on $r$ and can be approximated by ‘just adding one’ (Gates et al. 2020). This relation, valid for a Kerr black hole, can be rewritten for a Schwarzschild black hole as
$\frac{b}{M}\approx\frac{1}{2}\left[\frac{r}{M}+1+\sqrt{\left(\frac{r}{M}\right)^{2}+2\frac{r}{M}-1}\>\right]\approx\frac{r}{M}+1\hskip 6.99997pt\text{(face-on)}\>.$
@@
Hence, the event horizon, situated at the radius $r=2M$, appears at an apparent approximated radius $b\approx 2.82M$, in agreement with the numerical outputs of Figure 2, while outer radii develop towards higher values of $b$. As for geodesics of order $n=1$, if one stops the backward ray-tracing integration, properly introduced in Section 5, before getting the third crossing of the plane, the numerical inner and outer edges of the $n=1$ lensing band appear at $b=5.02M$ and $b=6.17M$ respectively (Gralla et al. 2019); these values agree with Figure 2.
@@
If the screen and the equatorial plane are not parallel, the images of equatorial rings are deformed by projection effects. [^8] In a Newtonian context, they would give rise to ellipses. On the contrary, the critical curve does not vary under change of the plane’s inclination, $i$, because of the spherical symmetry of the photon sphere and it always takes the shape of a circle of radius $b=3\sqrt{3}M$ on the plane of sky. [^9] In Kerr, the photon shell has an equatorial reflection symmetry and the critical curve is always symmetric with respect to the axis perpendicular to the projected direction of the spin (Gralla & Lupsasca 2020a).
@@
Next, we focused on the general properties of projections on the observer’s screen of equatorial circles as seen by an inclined observer. The discussions of the following paragraphs are complemented by additional details given in Appendix A. First of all, both the zeroth-order and the first-lensed projections of planar circular rings are equal to their non-inclined counterparts along the line on the screen parallel to the inclination’s axis of rotation of the plane (horizontal axis in the upper panels of Figure 1). Moreover, projected rings always possess a mirror symmetry on the observer’s screen with respect to the axis that is normal to the inclination’s one [^10] This property, reliant on the planar motion typical of all static spherical symmetric space-times, is broken in Kerr (Gralla & Lupsasca 2020a). (vertical axis in Figure 1).
@@
Let $\psi$ be the polar angle on the observer’s screen as shown in Figure 1. If the radius of the inclined ring satisfies $r_{\mathrm{H}}<r\leq\tilde{r}$, the zeroth-order projection of the side of the ring tilted backwards ($0^{\circ}<\psi<180^{\circ}$, ) or forwards ($180^{\circ}<\psi<360^{\circ}$) appears, respectively, at a higher or lower radial distance on the screen, $b$, with respect to the non-inclined apparent circle. As a result, since in our simulations the forward side of the plane points towards $\psi=270^{\circ}$, the impact parameter, $b$, of the inclined projected equatorial horizon, as well as that of rings with $r\leq 3M$, decreases from $\psi=90^{\circ}$ to $\psi=270^{\circ}$ (see grey and magenta lines in the upper panels of Figure 2). The same behaviour of $b$ is followed for the forward part of rings with $r>3M$: it then decreases from $\psi=180^{\circ}$ to $\psi=270^{\circ}$ in our set-up (see green and yellow lines in the upper panels of Figure 2). Instead, for the backward points of rings with $r>3M$, for a given pair of plane’s inclination and direction on the observer’s screen, there exist a radius for which the behaviour of $b$ is reversed, that is to say that the projection of a given point has a smaller impact parameter than its corresponding non-inclined projection. The existence of this radius, which is smaller for higher inclinations and for directions on the screen closer to the axis of reflection, is a necessary consequence of the flat space-time limit: in a Minkowski space-time, the projected image of a circle is an ellipse whose semi-major axis is along the rotation axis (the horizontal direction here) and equal to the non-inclined projection. Since the semi-minor axis is smaller than the semi-major axis by definition, a reversal must appear at large distance from the black hole to reproduce the Minkowski behaviour. For the small inclination considered in the set-up of Figure 2, this effect appears beyond the considered field of view.
@@
As for the first-lensed projection of equatorial rings, if the radius of the ring satisfies $r_{\mathrm{H}}<r\leq\tilde{r}$, the projection of the forward or backward side of the ring appears, respectively, at a higher or lower radial distance on the screen, $b$, with respect to the non-inclined apparent circle. Hence, as the geodesic executes a half turn around the black hole before touching the screen and since in our simulations the forward side of the plane points towards $\psi=270^{\circ}$, the impact parameter, $b$, of such rings decreases from $\psi=90^{\circ}$ to $\psi=270^{\circ}$ (see grey and magenta lines in the lower panels of Figure 2). If $r=3M$, there is no variation with respect to the non-inclined case (see green line in the lower panels of Figure 2). Finally, if $r>3M$, the modulations of $b$ are reversed. Thus, in our simulations, $b$ increases from $\psi=90^{\circ}$ to $\psi=270^{\circ}$ (see yellow line in the lower panels of Figure 2). Put differently, the ring projections flee the critical curve in the lower part of the screen and are attracted to it in the upper part.
2.4 Inner shadow, primary image, and photon rings
@@
The relevant observables appear on the far-away observer’s screen if and only if emitting material is present around the black hole. In particular, a single point source gives rise to an infinite sequence of images. These additional lensed observables were originally called ‘ghost images’ (Darwin 1959) and were present in the first simulated ‘photograph’ of an accretion disc around a Schwarzschild [^11] The first image of an accretion disc around a Kerr black hole with $a\neq 0$ has been computed by Viergutz (Viergutz 1993). black hole (Luminet 1979). Supermassive black holes such as Sgr A* and M87* are surrounded by emitting matter forming a hot, geometrically thick, and optically thin radiatively inefficient accretion flow (RIAF, Yuan & Narayan 2014). Moreover, for M87*, EHT data favours the scenario of a magnetically arrested disc (MAD) whose millimetre wavelength emission arises near the event horizon, predominantly in the vicinity of the equatorial plane (Narayan et al. 2003; The Event Horizon Telescope Collaboration 2019c; The Event Horizon Telescope Collaboration 2021).
@@
Images of black holes surrounded by optically thin emission display a brightness depression in their inner region, known as observable shadow, which does not always track the critical curve, but it is instead highly astrophysical dependent (Chael et al. 2021). For instance, for an idealistically thin equatorial disc, this ‘inner shadow’ can be well inside the critical curve: when the disc extends below the photon sphere, the edge of this darkest region corresponds to the projection of the inner radius of the equatorial disc modulated by redshift effects (Chael et al. 2021). For a disc seen face-on and extending to the event horizon of a Schwarzschild black hole, the inner shadow has a radius of about $2.82M$ on the observer’s screen, as discussed in Paragraph 2.3.
@@
The primary image is the one reaching the observer on geodesics touching the emitting equatorial source just once. It is also called $n=0$ or direct image. Photon rings are higher-order images of the same emitting material, labelled by the number $n\geq 1$ of the equatorial-plane crossings along their nearly bound orbit. When the source is intersected during the crossing, a flux increment is added. Also, the $n$-th photon ring of the equatorial source lies inside the $n$-th lensing band (Gralla & Lupsasca 2020a).
@@
When the emitting source is not isotropically distributed, as in the case of equatorial discs, photon rings exponentially tend to the critical curve, have exponentially decreasing angular width and thus are exponentially demagnified and exponentially loose flux density [^12] The first M87* photon ring provides $\sim 10\%$ of the total luminosity in numerical GRMHD simulations (Johnson et al. 2020). for increasing $n$ (Johnson et al. 2020; Vincent et al. 2022). Also, in intensity cross sections of the images, they do not blend with the direct image, unlike an infalling sphere, but they give birth to a stack of discrete higher-order images superimposed on top of the primary one (Gralla et al. 2019; Vincent et al. 2022), provided that absorption effects are weak at a given observing frequency (Beckwith & Done 2005). These properties are presented more meticulously in the introduction of Paugnat et al. (2022).
@@
To summarise, photon rings are signatures related to the strong-field lensing effects of the photon sphere and, unlike the critical curve or the lensing bands, they can be observed and possess only limited astrophysical dependence (Vincent et al. 2022). The present EHT does not resolve the photon rings (Gralla 2021), but their interferometric detection may soon become possible by observing at higher frequencies (ngEHT, Johnson et al. 2023) or through longer Earth-space baselines (BHEX, Johnson et al. 2024).
@@ 3.1 M87* distance and orientation
Our compact object model mimics the observationally assessed properties of M87*, namely its mass, $M$, its distance, $D$, and its inclination, $i$. The values of $M$ and $D$ are those inferred by the EHT collaboration (The Event Horizon Telescope Collaboration 2019d), the mass being consistent with that obtained via a stellar-dynamics study (Gebhardt et al. 2011). For its inclination, defined as the angle between the line of sight and the angular momentum of the accretion flow, we used the value of the viewing angle of the jet in M87 (Mertens et al. 2016), assuming that its axis is aligned with the angular momentum vector of the disc and that the latter points away from the observer.
@@ 3.2 Rezzolla-Zhidenko metric
In this paper, we consider the Rezzolla-Zhidenko metric (RZ metric, Rezzolla & Zhidenko 2014), which describes in a parametric framework the space-time outside a general spherically symmetric, [^13] The Konoplya-Rezzolla-Zhidenko metric (KRZ metric, Konoplya et al. 2016) generalises the RZ parametrisation to circular space-times, then to rotating black holes, but not all stationary axisymmetric black holes are included in this class. stationary, and therefore static black hole. [^14] Kocherlakota & Rezzolla (2020) extend the RZ scheme to arbitrary asymptotically flat, spherical symmetric, and static space-times, including non–black hole cases such as boson stars or naked singularities. In the strong-field regime, contrary to previous parametrisations in which an infinite number of equally important parameters is needed (Johannsen & Psaltis 2011), the RZ metric is described by a finite number of hierarchical factors. The line element of this metric in a spherical polar coordinate system $(t,r,\theta,\varphi)$ is
@@
where the functions $N$ and $B$ only depend on the radial coordinate, $r$. Conveniently for a black hole, $r$ can be compactified in the dimensionless variable $x\coloneqq 1-r_{\mathrm{H}}/r$, where $r_{\mathrm{H}}$ is the radial position of the event horizon so that $x=0$ corresponds to the location of the horizon and $x=1$ to spatial infinity. In terms of $x$, using $G=c=1$ natural units and introducing $\epsilon\coloneqq 2M/r_{\mathrm{H}}-1$, the functions $N$ and $B$ can be written as follows:
$\cases{N}^{2}(x)=x[1-\epsilon(1-x)+(a_{0}-\epsilon)(1-x)^{2}+\tilde{A}(x)(1-x)^{3}]\\ B(x)=1+b_{0}(1-x)+\tilde{B}(x)(1-x)^{2}{}\hskip-14.22636pt\>$, (5)
@@
$\tilde{A}(x)=\frac{a_{1}}{1+\dfrac{a_{2}\,x}{1+\dfrac{a_{3}\,x}{1+\ldots}}}\quad\color[rgb]{0,0,0},\quad\tilde{B}(x)=\frac{b_{1}}{1+\dfrac{b_{2}\,x}{1+\dfrac{b_{3}\,x}{1+\ldots}}}\>\>\,.$ (6)
The dimensionless constants $a_{0},a_{1},a_{2},a_{3}\ldots$ and $b_{0},b_{1},b_{2},b_{3}\ldots$ have to be constrained by observations. If all $\epsilon$, $a_{i}$, and $b_{i}$ are zero, with $i\in\mathbb{N}$, the Schwarzschild metric is recovered in Schwarzschild-Droste coordinates.
It is useful to notice that at the horizon the functions $\tilde{A}$, $\tilde{B}$ of equations (6), reduce to $\tilde{A}(0)=a_{1}$, $\tilde{B}(0)=b_{1}$. This means that near the horizon, only the lower-order terms in the expansion are relevant. Higher-order terms increase the accuracy of the approximation and, for numerous known metrics deviating slightly from the Schwarzschild one, the first order is sufficient to accurately investigate astrophysical observable deviations (Konoplya & Zhidenko 2020). Besides, the coefficients $a_{0}$ and $b_{0}$, which have been introduced in the equations (5), can be expressed as a combination of the parametrised post-Newtonian (PPN) parameters and, assuming black hole uniqueness, they are constrained by observations to be at most of the order of $10^{-4}$ (Will 2014).
For the sake of simplicity, as we were interested in the near-horizon emission around a black hole and following the arguments stated before, in the simulations all the parameters $a_{i}$, $b_{i}$ with $i\geq 2$ were put to zero. The parameters $a_{0}$ and $b_{0}$ were also fixed to be zero, their impact on the image formation being negligible with respect to that of the astrophysical set-up when PPN constraints are met, as it is illustrated in Appendix C; in the same Appendix we also present higher deviations in $a_{0}$ and $b_{0}$, allowed in theories of gravity without a Birkhoff-like uniqueness theorem. Finally, we chose $b_{1}=0$, again for simplicity purposes and because its variation has a very limited effect on the image, as shown in Section 6 and Appendix C. We could then study the effect on the images of the only unconstrained parameters, $a_{1}$ and $\epsilon$, which we made vary independently by setting the other parameter to zero. We chose them in their theoretically allowed [^15] The outermost horizon is located at $r=r_{\mathrm{H}}$ and no larger roots of $g_{rr}^{-1}=N^{2}(r)$ should exist (Kocherlakota & Rezzolla 2022). ranges and took ‘small’, in the sense of the following right inequality, [^16] This does not imply that the deviations are physically small. parametric deviations (Kocherlakota & Rezzolla 2020):
$1<\epsilon\leq 0.5,\quad a_{1}\geq-1\quad\text{and}\quad\max{(\,|\,a_{1}\,|,|\,\epsilon\,|\,)}\leq 1\>.$
Table 1: Parameters of the black hole.